# Post-Quantum Cryptography (PQC) Network Instrument: Measuring PQC Adoption Rates and Identifying Migration Pathways

### v2

@@ I Introduction
The problem of adopting quantum-resistant cryptographic network protocols or post-quantum cryptography (PQC) is critically important to democratizing quantum computing. The problem is urgent because practical quantum computers [1], will break classical encryption in the next few decades. Major applications such as financial services, high-performance computing (HPC) or supercomputing, and health analytics must be migrated to be quantum-resistant. The main challenges of adopting post-quantum cryptography lie in both algorithmic complexity and hardware/software/network implementation. The grand question of how existing cyberinfrastructure will support post-quantum cryptography remains unanswered.
This paper describes: i) the design of a Post-Quantum Cryptography (PQC) network instrument placed in a national-scale supercomputing center, ii) the latest results on PQC adoption rate across a wide spectrum of popular network protocols (Secure Shell – SSH, Transport Layer Security – TLS, etc.), iii) the current state of PQC implementation in key scientific applications (e.g., OpenSSH, SciTokens), iv) the challenges of being quantum-resistant and v) discussion of potential novel attacks. Our result is critically important regarding adopting National Institute of Standards and Technology (NIST)’s draft algorithms, such as CRYSTALS-Kyber for encryption, FALCON and SPINCS+ for digital signature, and KEMTALS for key exchange [2] and protocol-specific PQC adaptation such as Hybrid Streamlined NTRU (Ring-Based Public Key Cryptosystem) Prime sntrup761 and x25519 with SHA-512 (sntrup761x25519-sha512) in OpenSSH [3, 4].
@@
With new developments in quantum computers and quantum algorithms on the rise, many of these problems previously assumed to be hard appear much easier. The 1994 Shor’s algorithm for solving both discrete logs and prime factorization efficiently on quantum computers shows the need for this Post-Quantum Cryptography (PQC), which is made even more urgent when you consider an adversary who is storing classically encrypted traffic now to be decrypted with a quantum computer in the near future when they are powerful and reliable enough. The grand question of how existing cyberinfrastructure will support post-quantum cryptography remains unanswered and will be addressed in this paper.
@@
Our analyses identify pathways to migrate current applications to be quantum-resistant. Through experimental deployment at different network vantage points, our approach may help identify empirical novel attacks against PQC implementations and give feedback to the National Institute of Standards and Technology (NIST) as well as developers to improve their drafts and successfully migrate to PQC [7].
@@
Remark 1: Perspective w.r.t the state of the art. To the best of our knowledge, this is the first network instrument that exists to monitor PQC adoption publicly. There are a few related work [9, 10, 11, 12], e.g., Cloudflare released a snapshot of TLS adoption in Feb 2024 through their content delivery network [10]. On the other hand, our network instrument is the first of such to monitor the adoption of both OpenSSH, TLS, and other protocols at large-scale, high-speed (gigabit to terabit) interconnected networks with a wide spectrum of not only traditional cloud but also scientific workloads.
@@ II Approach Overview
Ethical considerations. Our approach works on connection metadata produced by Zeek. We did not use man-in-the-middle or traditional SSL termination equipment. Thus, we only see the encrypted data and not the original content. However, we do have access to the timing information of connection establishment to derive the trend and seasonality of the PQC adoption rate. The connection metadata generally includes standard information such as IP address, certificate, and negotiation of cipher suites. We do not see Personally Identifiable Information (PII).
@@
Remark 2: Our Zeek-integrated network instrument to measure PQC adoption has three main components: Parallel log parser to ingest large-volume of Zeek logs (metadata) from a +400Gbps network. Historical analyses scripts that perform regression, trend, and statistics on the parsed metadata of cipher suites. Real-time snapshot and comparison of our results vs. others (if any) to detect novel attacks and give feedback to the National Institute of Standards and Technology (NIST).
@@ III Datasets
Remark 3: Our dataset is unique in its placement at a strategic network vantage point to observe metadata of a wide spectrum of applications across different protocols, validating our PQC measurement for both historical and real-time adoption rates.
@@ IV-A Background on PQC adoption in OpenSSH
Remark 4: The sntrup761 KEM is a lattice-based post-quantum algorithm. Lattice-based algorithms are relatively new in the cryptography world and are thus not as strongly tested against even pre-quantum adversaries as some more traditional schemes. Many PQC implementations such as this one thus combine a more established cryptosystem with the intention of ensuring immediate security if just one of the involved algorithms are proven weak [29].
@@ IV-B Results on PQC protocols adoption in SSH
Remark 5: SSH is among the first protocols that implement PQC and actually have clients adopt their PQC protocol widely, as shown in our results.
@@ IV-C Results on classical protocols in SSH
Through further inspection of metadata logs, we determined that over 83% of Server-side SSH protocols were from 2019 and earlier, three years before sntrup761x25519 was even introduced. This highlights a common problem in the technology space—the difficulty of ensuring all systems are updated and using the most secure connection options available. This is important if we are to ever reach full PQC adoption.
@@ V Method and Results
This section describes our analysis extended to other major network protocols such as Remote Desktop Protocol (RDP), Domain Name System (DNS), and Transport Layer Security (TLS). Note that this analysis is focused on a sample working day to illustrate the adoption. The objective is to measure the readiness level of applications using these protocols with regard to Post-Quantum Cryptography.
@@ V-A RDP
Even though for RDP, there was a minimal data set of 26 connections, we were able to identify an interesting issue. Generally, RDP can be configured to use two types of cryptographic security, Enhanced Encryption and Network-Layer Authentication (NLA). Enhanced Encryption for RDP offers encryption over TLS of everything after the Connection Initiation stage and also allows authentication of the server from the client side. NLA requires the client to be authenticated before establishing a session with the server. However, out of all 26 connections, only 2 used the HYBRID-EX security protocol setting, which uses both Enhanced Encryption and NLA. As with SSH, this calls for more careful management of individual machines and making sure they are all configured to use both Enhanced Encryption and NLA for optimal security.
@@ V-C TLS
Remark 6: This difficulty in adopting TLS v1.3, though, is still prevalent throughout the internet and not just here at NCSA, and it brings to light the unfortunate truth that if widely adopting the newest TLS has taken this many years, PQC adoption may show the same challenge.
@@ VI Current PQC adoption of network protocols and their applications and libraries
Remark 7: The key takeaway from our study is that only SSH and SSL have had PQC implementation and some real-world adoptions. The rest of the protocols have to roll their own crypto or encapsulate their data in TLS 1.3 (hybrid quantum key exchange).
@@ VII Case study: Challenges of Migrating
Background. The SciTokens protocol uses JSON web tokens (JWTs) between various parties, which will validate any access to resources. The process starts off by including information within the JWT. The information included most commonly consists of who issued the token, who the token is assigned to, when the token will expire, and what permissions and access the end user of such token will have. The tokens are then signed cryptographically to be authentic and certified when called upon to use. The benefit of this approach is that each organization or institution can issue its own tokens, allowing for easier and quicker access to resources without relying on one central authority to control the distribution of such tokens.
Making SciTokens Quantum Resistant. To be quantum-resistant, existing software must migrate encrypting in three main steps: signing, key exchange, and encryption. Key exchange and encryption are currently effective, but the looming threat of quantum computing could compromise these safeguards, potentially exposing sensitive data. For SciTokens, the main critical component that must be quantum secure is the signing step. When a digital signature is forged, the root of trust about the integrity of the signed token is broken. This undermines trust in historical transaction records, as the digital signatures securing them could be compromised by a quantum computer.
@@
Remark 8: To the developer community, it is unclear what is the process of determining which post-quantum secure algorithm to use and how to implement those correctly. In the future, we will work with the SciTokens community to correctly implement its quantum-resistant digital signature algorithm.
@@ VIII Lessons Learned &
Preempting PQC attacks. Despite numerous attack attempts  [47] and analyses against current PQC drafts, none of the attacks have been publicly confirmed. We assert that if there were any successful attacks, they would highly likely leave traces in the network metadata, which would be measured by our network instrument or a honeypot [48], given that we can achieve a widespread deployment on the Internet scale. Therefore, it is important to continue this line of network instruments for PQC measurement to preempt attacks [49] and failures [50], particularly in HPC environments.
@@
Remark 9: We are releasing an open dataset containing metadata of cryptographic suites across major network protocols (SSH, TLS, RDP, etc.) upon the paper’s publication at https://pmcao.github.io/pqc. With our network instrument in place, we will track the adoption of PQC algorithms, give feedback to NIST, expect to discover potential novel PQC attacks in the wild, and help migrate HPC applications to be quantum-safe. Finally, we will expand our network instrument to a set of networked vantage points (PQC telescopes) to continuously measure PQC adoption at the Internet scale.

### v3

@@ I Introduction
The problem of adopting quantum-resistant cryptographic network protocols or post-quantum cryptography (PQC) is critically important to democratizing quantum computing. The problem is urgent because practical quantum computers [1], will break classical encryption in the next few decades. Major applications such as cloud computing [2] (including HPC/supercomputing), financial services, and health analytics [3] must be migrated to be quantum-resistant. The main challenges of adopting post-quantum cryptography lie in algorithmic complexity and hardware/software/network implementation. The grand question of how existing cyberinfrastructure will support post-quantum cryptography remains unanswered.
This paper describes: i) the design of a Post-Quantum Cryptography (PQC) network instrument placed in a national-scale supercomputing center, ii) the latest results on PQC adoption rate across a wide spectrum of widely used network protocols (Secure Shell – SSH, Transport Layer Security – TLS, etc.), iii) the current state of PQC implementation in key scientific applications (e.g., OpenSSH, SciTokens), iv) the challenges of being quantum-resistant and v) discussion of potential novel attacks. Our result is critically important regarding adopting National Institute of Standards and Technology (NIST)’s draft algorithms, such as CRYSTALS-Kyber for encryption, FALCON and SPINCS+ for digital signature, and KEMTALS for key exchange [4] and protocol-specific PQC adaptation such as Hybrid Streamlined NTRU (Ring-Based Public Key Cryptosystem) Prime sntrup761 and x25519 with SHA-512 (sntrup761x25519-sha512) in OpenSSH [5, 6].
@@
With new developments in quantum computers and quantum algorithms on the rise, many of these problems previously assumed to be hard appear much easier. The 1994 Shor’s algorithm for solving both discrete logs and prime factorization efficiently on quantum computers shows the need for this Post-Quantum Cryptography (PQC), which is made even more urgent when you consider an adversary who is storing classically encrypted traffic now to be decrypted with a quantum computer soon when they are powerful and reliable enough. The grand question of how existing cyberinfrastructure will support post-quantum cryptography remains unanswered and will be addressed in this paper.
@@
Remark 1: Our research vs. the state of the art. To the best of our knowledge, this is the first network instrument that exists to monitor PQC adoption publicly. There are a few related work [9, 10, 11, 12], e.g., Cloudflare released a snapshot of TLS adoption in Feb 2024 through their content delivery network [10]. On the other hand, our network instrument is the first of such to monitor the adoption of both OpenSSH, TLS, and other protocols at large-scale, high-speed (gigabit to terabit) interconnected networks with a wide spectrum of not only traditional cloud but also scientific workloads.
Remark 2: Our instrument is tightly integrated with Zeek and network border gateways. Our Zeek-integrated network instrument to measure PQC adoption has three main components: Parallel and memory-safe log parser to ingest large-volume Zeek logs from a TeraCore network. Historical analyses scripts that perform regression, trend, and statistics on the parsed metadata of cipher suites. Real-time snapshot and comparison of our results vs. others (if any) to detect novel attacks and give feedback to NIST and security operators.
Through experimental deployment at the edges, our approach may help identify empirical novel attacks against PQC implementations and give feedback to the National Institute of Standards and Technology (NIST) and migration pathways for HPC developers [13].
@@ II Approach Overview
Ethical considerations. Our approach works on connection metadata produced by Zeek. We did not use man-in-the-middle or traditional SSL termination equipment. Thus, we only see the encrypted data and not the original content. However, we have access to the timing information of connection establishment to derive the trend and seasonality of the PQC adoption rate. The connection metadata generally includes standard information such as IP address, certificate, and negotiation of cipher suites. We do not see Personally Identifiable Information (PII).
Remark 3: Uniqueness of our dataset. Our dataset is unique in its placement at a strategic network vantage point at the border gateway router to observe metadata of a wide spectrum of applications across different protocols, validating our PQC measurement for both historical and real-time adoption rates.
@@ IV-A Background on PQC adoption in OpenSSH
Remark 4: Quantum-resistant property of Streamlined NTRU Prime sntrup761. The sntrup761 KEM is a lattice-based post-quantum algorithm. Lattice-based algorithms are relatively new in the cryptography world and are thus not as strongly tested against even pre-quantum adversaries as some more traditional schemes. Many PQC implementations such as this one thus combine a more established cryptosystem to ensure immediate security if just one of the involved algorithms is proven weak [31].
@@ IV-B Results on PQC protocols adoption in SSH
Remark 5: Secure Shell (SSH) are among first widely adopted PQC implementation. SSH is among the first protocols to implement PQC, albeit with a low adoption rate of 0.029% recorded at NCSA. However, as our results show, clients increasingly adopt PQC in SSH.
@@ IV-C Results on classical protocols in SSH
Through further inspection of metadata logs, we determined that over 83% of Server-side SSH protocols were from 2019 and earlier, three years before sntrup761x25519 was even introduced. This highlights a common problem in the technology space—the difficulty of ensuring all systems are updated and using the most secure connection options available. This is important if we are to ever reach a complete PQC adoption.
@@ V Method and Results
This section describes our analysis extended to other major network protocols such as Remote Desktop Protocol (RDP), Domain Name System (DNS), and Transport Layer Security (TLS). Note that this analysis is focused on a sample working day to illustrate the adoption. The objective is to measure the readiness level of applications using these protocols with regard to Post-Quantum Cryptography. As we already provided the trend and seasonality of SSH adoption over the last year, in this part, we will provide a snapshot of other protocols in a working day of collected network traffic.
@@ V-A RDP
Even though for RDP, there was a minimal data set of 26 connections, we were able to identify an interesting issue. Generally, RDP can be configured to use two types of cryptographic security, Enhanced Encryption and Network-Layer Authentication (NLA). Enhanced Encryption for RDP offers encryption over TLS of everything after the Connection Initiation stage and also allows authentication of the server from the client side. NLA requires the client to be authenticated before establishing a session with the server. However, out of all 26 connections, only two of them used the HYBRID-EX security protocol setting, which uses both Enhanced Encryption and NLA. As with SSH, this calls for more careful management of individual machines and making sure they are all configured to use both Enhanced Encryption and NLA for optimal security.
@@
Remark 6: Initial adoption of TLS v1.3 with hybrid PQC protocol is relatively slow. This difficulty in adopting TLS v1.3, though, is still prevalent throughout the internet and not just here at NCSA. It brings to light the unfortunate truth that if widely adopting the newest TLS has taken this many years, PQC adoption may show the same challenge.
@@ VI Current PQC adoption of network protocols and their applications and libraries
Remark 7: Other than SSH and TLS, other network protocols are not ready for PQC. Only SSH and SSL have had PQC implementation with real-world adoptions. The rest of the protocols must develop their quantum-resistant cryptographic system or encapsulate their data in TLS 1.3 (hybrid quantum key exchange).
@@ VII Case study: Challenges of Migrating
Background. The SciTokens protocol uses JSON web tokens (JWTs) between various parties, which will validate any access to resources. The process starts off by including information within the JWT. The information included most commonly consists of who issued the token, who the token is assigned to, when the token will expire, and what permissions and access the end user of such token will have. The tokens are then signed cryptographically to be authentic and certified when called upon to use. The benefit of this approach is that each organization or institution can issue its tokens, allowing for easier and quicker access to resources without relying on one central authority to control the distribution of such tokens.
Securing SciTokens to be Quantum Resistant. Existing efforts to secure token-based authentications include formal verification [43]. The next level of security guarantee is to make token-based authentication quantum-resistant. For this, existing software must migrate cryptographic protocols in three main steps: signing, key exchange, and encryption. Key exchange and encryption are currently effective, but the looming threat of quantum computing could compromise these safeguards, potentially exposing sensitive data. For SciTokens, the main critical component that must be quantum secure is the signing step. When a digital signature is forged, the root of trust about the integrity of the signed token is broken. This undermines trust in historical transaction records, as the digital signatures securing them could be compromised by a quantum computer.
@@
Remark 8: Uncertainty in adopting PQC To the developer community, it is unclear what is the process of determining which post-quantum secure algorithm to use and how to implement those correctly. In the future, we will work with the SciTokens community to correctly implement its quantum-resistant digital signature algorithm.
@@ VIII Lessons Learned &
Observing and Preempting PQC attacks. Despite numerous attack attempts  [50] and analyses against current PQC drafts, none of the attacks have been publicly confirmed. We assert that if there were any successful attacks, they would highly likely leave traces in the network metadata, which would be measured by our network instrument or a honeypot [51], given that we can achieve a widespread deployment on the Internet scale. Therefore, it is important to continue this line of network instruments for PQC measurement to preempt attacks [52, 53] and failures [54], particularly in HPC environments. We suspect that futuristic malware [55] may employ PQC, in addition to other techniques such as Machine Learning, to hide themselves from forensic analyses.
@@
Remark 9: An open dataset containing metadata of cryptographic suites for major network protocols. We are releasing an open dataset containing metadata of cryptographic suites across major network protocols (SSH, TLS, RDP, etc.) upon the paper’s publication at https://pmcao.github.io/pqc. With our network instrument in place, we will track the adoption of PQC algorithms, give feedback to NIST, expect to discover potential novel PQC attacks in the wild, and help migrate HPC applications to be quantum-safe. Finally, we will expand our network instrument to a set of networked vantage points (PQC telescopes) to continuously measure PQC adoption at the Internet scale.